\begin{thebibliography}{10}
\expandafter\ifx\csname url\endcsname\relax
  \def\url#1{\texttt{#1}}\fi
\expandafter\ifx\csname urlprefix\endcsname\relax\def\urlprefix{URL }\fi
\expandafter\ifx\csname href\endcsname\relax
  \def\href#1#2{#2} \def\path#1{#1}\fi

\bibitem{Lyapunov-1892}
A.~Lyapunov, The General Problem of the Stability of Motion (in Russian),
  Kharkov, 1892, [English transl.: Academic Press, NY, 1966].

\bibitem{GrassbergerP-1983}
P.~Grassberger, I.~Procaccia, Measuring the strangeness of strange attractors,
  Physica D: Nonlinear Phenomena 9~(1-2) (1983) 189--208.

\bibitem{EckmannR-1985}
J.-P. Eckmann, D.~Ruelle, Ergodic theory of chaos and strange attractors,
  Reviews of Modern Physics 57~(3) (1985) 617--656.

\bibitem{ConstantinFT-1985}
P.~Constantin, C.~Foias, R.~Temam, Attractors representing turbulent flows,
  Memoirs of the American Mathematical Society 53~(314) (1985) 1--67.

\bibitem{AbarbanelBST-1993}
H.~Abarbanel, R.~Brown, J.~Sidorowich, L.~Tsimring, The analysis of observed
  chaotic data in physical systems, Reviews of Modern Physics 65~(4) (1993)
  1331--1392.

\bibitem{KaplanY-1979}
J.~Kaplan, J.~Yorke, Chaotic behavior of multidimensional difference equations,
  in: Functional Differential Equations and Approximations of Fixed Points,
  Springer, Berlin, 1979, pp. 204--227.

\bibitem{Kolmogorov-1959}
A.~Kolmogorov, On entropy per unit time as a metric invariant of automorphisms,
  Dokl. Akad. Nauk SSSR (In Russian) 124~(4) (1959) 754--755.

\bibitem{Sinai-1959}
Y.~Sinai, On the notion of entropy of dynamical systems, Dokl. Akad. Nauk SSSR
  (In Russian) 124~(4) (1959) 768--771.

\bibitem{AdlerKA-1965}
R.~Adler, A.~Konheim, M.~McAndrew, Topological entropy, Transactions of the
  American Mathematical Society 114~(2) (1965) 309?319.

\bibitem{Millionschikov-1976}
V.~M. Millionschikov, A formula for the entropy of smooth dynamical systems,
  Differencial'nye Uravenija (in Russian) 12~(12) (1976) 2188--2192, 2300.

\bibitem{Pesin-1977}
Y.~Pesin, Characteristic {L}yapunov exponents and smooth ergodic theory,
  Russian Mathematical Surveys 32~(4) (1977) 55--114.

\bibitem{Young-2013}
L.-S. Young, Mathematical theory of {L}yapunov exponents, Journal of Physics A:
  Mathematical and Theoretical 46~(25) (2013) 254001.

\bibitem{CorintoF-2017}
F.~Corinto, M.~Forti, Memristor circuits: bifurcations without parameters, IEEE
  Transactions on Circuits and Systems I: Regular Papers 64~(6) (2017)
  1540--1551.

\bibitem{Tetzlaff-2014-book}
R.~Tetzlaff (Ed.), Memristors and Memristive Systems, Springer, 2014.

\bibitem{Chua-1993}
L.~Chua, Global unfolding of {C}hua's circuit, IEICE Transactions on
  Fundamentals of Electronics, Communications, Computer Science E76-A (1993)
  704--734.

\bibitem{HuangPWF-1996}
A.~Huang, L.~Pivka, C.~Wu, M.~Franz, Chua's equation with cubic nonlinearity,
  International Journal of Bifurcation and Chaos 06~(12a) (1996) 2175--2222.

\bibitem{BilottaP-2008}
E.~Bilotta, P.~Pantano, A gallery of {C}hua attractors, Vol. Series~A. 61,
  World Scientific, 2008.

\bibitem{KuznetsovLV-2010-IFAC}
N.~Kuznetsov, G.~Leonov, V.~Vagaitsev, Analytical-numerical method for
  attractor localization of generalized {C}hua's system, IFAC Proceedings
  Volumes 43~(11) (2010) 29--33.
\newblock \href {http://dx.doi.org/10.3182/20100826-3-TR-4016.00009}
  {\path{doi:10.3182/20100826-3-TR-4016.00009}}.

\bibitem{LeonovKV-2011-PLA}
G.~Leonov, N.~Kuznetsov, V.~Vagaitsev, Localization of hidden {C}hua's
  attractors, Physics Letters A 375~(23) (2011) 2230--2233.
\newblock \href {http://dx.doi.org/10.1016/j.physleta.2011.04.037}
  {\path{doi:10.1016/j.physleta.2011.04.037}}.

\bibitem{BraginVKL-2011}
V.~Bragin, V.~Vagaitsev, N.~Kuznetsov, G.~Leonov, Algorithms for finding hidden
  oscillations in nonlinear systems. {T}he {A}izerman and {K}alman conjectures
  and {C}hua's circuits, Journal of Computer and Systems Sciences International
  50~(4) (2011) 511--543.
\newblock \href {http://dx.doi.org/10.1134/S106423071104006X}
  {\path{doi:10.1134/S106423071104006X}}.

\bibitem{LeonovKV-2012-PhysD}
G.~Leonov, N.~Kuznetsov, V.~Vagaitsev, Hidden attractor in smooth {C}hua
  systems, Physica D: Nonlinear Phenomena 241~(18) (2012) 1482--1486.
\newblock \href {http://dx.doi.org/10.1016/j.physd.2012.05.016}
  {\path{doi:10.1016/j.physd.2012.05.016}}.

\bibitem{KuznetsovKLV-2013}
N.~Kuznetsov, O.~Kuznetsova, G.~Leonov, V.~Vagaitsev, Analytical-numerical
  localization of hidden attractor in electrical {C}hua's circuit, Lecture
  Notes in Electrical Engineering 174~(4) (2013) 149--158.
\newblock \href {http://dx.doi.org/10.1007/978-3-642-31353-0\_11}
  {\path{doi:10.1007/978-3-642-31353-0\_11}}.

\bibitem{KiselevaKKKLYY-2017}
M.~Kiseleva, E.~Kudryashova, N.~Kuznetsov, O.~Kuznetsova, G.~Leonov,
  M.~Yuldashev, R.~Yuldashev, Hidden and self-excited attractors in {C}hua
  circuit: synchronization and {SPICE} simulation, International Journal of
  Parallel, Emergent and Distributed Systems (2017) 1--11\
  {d}oi:10.1080/17445760.2017.1334776.

\bibitem{StankevichKLC-2017}
N.~Stankevich, N.~Kuznetsov, G.~Leonov, L.~Chua, Scenario of the birth of
  hidden attractors in the {C}hua circuit, International Journal of Bifurcation
  and Chaos 27~(12), {a}rt. num. 1730038.

\bibitem{LeonovK-2013-IJBC}
G.~Leonov, N.~Kuznetsov, Hidden attractors in dynamical systems. {F}rom hidden
  oscillations in {H}ilbert-{K}olmogorov, {A}izerman, and {K}alman problems to
  hidden chaotic attractors in {C}hua circuits, International Journal of
  Bifurcation and Chaos 23~(1), {a}rt. no. 1330002.
\newblock \href {http://dx.doi.org/10.1142/S0218127413300024}
  {\path{doi:10.1142/S0218127413300024}}.

\bibitem{LeonovKM-2015-EPJST}
G.~Leonov, N.~Kuznetsov, T.~Mokaev, Homoclinic orbits, and self-excited and
  hidden attractors in a {L}orenz-like system describing convective fluid
  motion, Eur. Phys. J. Special Topics 224~(8) (2015) 1421--1458.
\newblock \href {http://dx.doi.org/10.1140/epjst/e2015-02470-3}
  {\path{doi:10.1140/epjst/e2015-02470-3}}.

\bibitem{Kuznetsov-2016}
N.~Kuznetsov, Hidden attractors in fundamental problems and engineering models.
  {A} short survey, Lecture Notes in Electrical Engineering 371 (2016) 13--25,
  (Plenary lecture at International Conference on Advanced Engineering Theory
  and Applications 2015).
\newblock \href {http://dx.doi.org/10.1007/978-3-319-27247-4\_2}
  {\path{doi:10.1007/978-3-319-27247-4\_2}}.

\bibitem{PhamVVJKH-2016}
V.-T. Pham, S.~Vaidyanathan, C.~Volos, S.~Jafari, N.~Kuznetsov, T.~Hoang, A
  novel memristive time--delay chaotic system without equilibrium points, The
  European Physical Journal Special Topics 225~(1) (2016) 127--136.

\bibitem{HuLLYZ-2017-HA}
X.~Hu, C.~Liu, L.~Liu, Y.~Yao, G.~Zheng, Multi-scroll hidden attractors and
  multi-wing hidden attractors in a 5-dimensional memristive system, Chinese
  Physics B 26~(11) (2017) 110502.

\bibitem{BaoBWCX-2017-HA}
B.~Bao, H.~Bao, N.~Wang, M.~Chen, Q.~Xu, Hidden extreme multistability in
  memristive hyperchaotic system, Chaos, Solitons \& Fractals 94 (2017) 102 --
  111.

\bibitem{BaoBLWW-2016-HA}
H.~Bao, B.~Bao, Y.~Lin, J.~Wang, H.-G. Wu, Hidden attractor and its dynamical
  characteristic in memristive self-oscillating system, Acta Physica Sinica 65,
  {art. num.} 180501.

\bibitem{PhamVVHV-2016-HA}
V.-T. Pham, S.~Vaidyanathan, C.~Volos, T.~Hoang, Y.~Van, Advances in Chaos
  Theory and Intelligent Control, Springer, 2016, Ch. Dynamics, synchronization
  and {SPICE} implementation of a memristive system with hidden hyperchaotic
  attractor, pp. 35--52.

\bibitem{Vaidyanathan-2017-SCI-HA}
S.~Vaidyanathan, A Memristor-Based Hyperchaotic System with Hidden Attractor
  and Its Sliding Mode Control, Springer International Publishing, 2017, pp.
  343--369.
\newblock \href {http://dx.doi.org/10.1007/978-3-319-55598-0_15}
  {\path{doi:10.1007/978-3-319-55598-0_15}}.

\bibitem{VaidyanathanPV-2017-HA}
S.~Vaidyanathan, V.-T. Pham, C.~Volos, Advances in Memristors, Memristive
  Devices and Systems, Springer, 2017, Ch. Adaptive Control, Synchronization
  and Circuit Simulation of a Memristor-Based Hyperchaotic System With Hidden
  Attractors, pp. 101--130.

\bibitem{PhamVVTT-2017-HA}
V.-T. Pham, S.~Vaidyanathan, C.~Volos, E.~Tlelo-Cuautle, F.~Tahir, A Memristive
  System with Hidden Attractors and Its Engineering Application, Springer
  International Publishing, 2017, pp. 81--99.

\bibitem{PhamVVWH-2017-HA}
V.-T. Pham, S.~Vaidyanathan, C.~Volos, X.~Wang, D.~Hoang, A Hyperjerk
  Memristive System with Hidden Attractors, Springer International Publishing,
  2017, pp. 59--80.

\bibitem{RochaRK-2017-HA}
R.~Rocha, J.~Ruthiramoorthy, T.~Kathamuthu, Memristive oscillator based on
  {C}hua's circuit: stability analysis and hidden dynamics, Nonlinear Dynamics
  88~(4) (2017) 2577–2587.

\bibitem{SahaSRC-2015-HA}
P.~Saha, D.~Saha, A.~Ray, A.~Chowdhury, Memristive non-linear system and hidden
  attractor, European Physical Journal: Special Topics 224~(8) (2015)
  1563--1574.

\bibitem{SemenovKASVA-2015}
V.~Semenov, I.~Korneev, P.~Arinushkin, G.~Strelkova, T.~Vadivasova,
  V.~Anishchenko, Numerical and experimental studies of attractors in
  memristor-based {C}hua's oscillator with a line of equilibria.
  {N}oise-induced effects, European Physical Journal: Special Topics 224~(8)
  (2015) 1553--1561.

\bibitem{PhamVJWV-2014-HA}
V.-T. Pham, C.~Volos, S.~Jafari, X.~Wang, S.~Vaidyanathan, Hidden hyperchaotic
  attractor in a novel simple memristive neural network, Optoelectronics and
  Advanced Materials - Rapid Communications 8~(11-12) (2014) 1157--1163.

\bibitem{PhamVVLV-2015-HA}
V.-T. Pham, C.~Volos, S.~Vaidyanathan, T.~Le, V.~Vu, A memristor-based
  hyperchaotic system with hidden attractors: Dynamics, synchronization and
  circuital emulating, Journal of Engineering Science and Technology Review 2
  (2015) 205--214.

\bibitem{ChenYB-2015-HA-EL}
M.~Chen, J.~Yu, B.-C. Bao, Finding hidden attractors in improved
  memristor-based {C}hua's circuit, Electronics Letters 51 (2015) 462--464.

\bibitem{ChenLYBXW-2015-HA}
M.~Chen, M.~Li, Q.~Yu, B.~Bao, Q.~Xu, J.~Wang, Dynamics of self-excited
  attractors and hidden attractors in generalized memristor-based {C}hua's
  circuit, Nonlinear Dynamics 81 (2015) 215--226.

\bibitem{AfraimovichGLShT-2014}
V.~S. {Afraimovich}, S.~V. {Gonchenko}, L.~M. {Lerman}, A.~L. {Shilnikov},
  D.~V. {Turaev}, Scientific heritage of {LP} {S}hilnikov, Regular and Chaotic
  Dynamics 19~(4) (2014) 435--460.

\bibitem{BianchiKLYY-2015}
G.~Bianchi, N.~Kuznetsov, G.~Leonov, M.~Yuldashev, R.~Yuldashev, Limitations of
  {PLL} simulation: hidden oscillations in {MATLAB} and {SPICE}, International
  Congress on Ultra Modern Telecommunications and Control Systems and Workshops
  (ICUMT 2015) 2016-January (2016) 79--84.
\newblock \href {http://dx.doi.org/10.1109/ICUMT.2015.7382409}
  {\path{doi:10.1109/ICUMT.2015.7382409}}.

\bibitem{BlagovKLYY-2015}
M.~Blagov, N.~Kuznetsov, G.~Leonov, M.~Yuldashev, R.~Yuldashev, Simulation of
  {PLL} with impulse signals in {MATLAB}: Limitations, hidden oscillations, and
  pull-in range, International Congress on Ultra Modern Telecommunications and
  Control Systems and Workshops (ICUMT 2015) 2016-January (2016) 85--90.
\newblock \href {http://dx.doi.org/10.1109/ICUMT.2015.7382410}
  {\path{doi:10.1109/ICUMT.2015.7382410}}.

\bibitem{LeonovKYY-2015-TCAS}
G.~Leonov, N.~Kuznetsov, M.~Yuldashev, R.~Yuldashev, Hold-in, pull-in, and
  lock-in ranges of {PLL} circuits: rigorous mathematical definitions and
  limitations of classical theory, IEEE Transactions on Circuits and
  Systems--I: Regular Papers 62~(10) (2015) 2454--2464.
\newblock \href {http://dx.doi.org/10.1109/TCSI.2015.2476295}
  {\path{doi:10.1109/TCSI.2015.2476295}}.

\bibitem{KuznetsovLYY-2017-CNSNS}
N.~Kuznetsov, G.~Leonov, M.~Yuldashev, R.~Yuldashev, Hidden attractors in
  dynamical models of phase-locked loop circuits: limitations of simulation in
  {MATLAB} and {SPICE}, Commun Nonlinear Sci Numer Simulat 51 (2017) 39--49.
\newblock \href {http://dx.doi.org/10.1016/j.cnsns.2017.03.010}
  {\path{doi:10.1016/j.cnsns.2017.03.010}}.

\bibitem{Kuznetsov-2016-PLA}
N.~Kuznetsov, The {L}yapunov dimension and its estimation via the {L}eonov
  method, Physics Letters A 380~(25--26) (2016) 2142--2149.
\newblock \href {http://dx.doi.org/10.1016/j.physleta.2016.04.036}
  {\path{doi:10.1016/j.physleta.2016.04.036}}.

\bibitem{KuznetsovLMPS-2017}
N.~Kuznetsov, G.~Leonov, T.~Mokaev, A.~Prasad, M.~Shrimali, {Finite-time
  Lyapunov dimension and hidden attractor of the Rabinovich system}, Nonlinear
  dynamics\ accepted, \url{https://arxiv.org/pdf/1504.04723}.

\bibitem{DouadyO-1980}
A.~Douady, J.~Oesterle, Dimension de {H}ausdorff des attracteurs, C.R. Acad.
  Sci. Paris, Ser. A. (in French) 290~(24) (1980) 1135--1138.

\bibitem{DoeringGHN-1987}
C.~Doering, J.~Gibbon, D.~Holm, B.~Nicolaenko, Exact {L}yapunov dimension of
  the universal attractor for the complex {G}inzburg-{L}andau equation, Phys.
  Rev. Lett. 59 (1987) 2911--2914.

\bibitem{KuznetsovAL-2016}
N.~Kuznetsov, T.~Alexeeva, G.~Leonov, Invariance of {L}yapunov exponents and
  {L}yapunov dimension for regular and irregular linearizations, Nonlinear
  Dynamics 85~(1) (2016) 195--201.
\newblock \href {http://dx.doi.org/10.1007/s11071-016-2678-4}
  {\path{doi:10.1007/s11071-016-2678-4}}.

\bibitem{Leonov-1991-Vest}
G.~Leonov, On estimations of {H}ausdorff dimension of attractors, Vestnik St.
  Petersburg University: Mathematics 24~(3) (1991) 38--41, [Transl. from
  Russian: Vestnik Leningradskogo Universiteta. Mathematika, 24(3), 1991,
  pp.~41-44].

\bibitem{LeonovB-1992}
G.~Leonov, V.~Boichenko, Lyapunov's direct method in the estimation of the
  {H}ausdorff dimension of attractors, Acta Applicandae Mathematicae 26~(1)
  (1992) 1--60.

\bibitem{Smith-1986}
R.~Smith, Some application of {H}ausdorff dimension inequalities for ordinary
  differential equation, Proc. Royal Society Edinburg 104A (1986) 235--259.

\bibitem{Leonov-2012-PMM}
G.~Leonov, Lyapunov functions in the attractors dimension theory, Journal of
  Applied Mathematics and Mechanics 76~(2) (2012) 129--141.

\bibitem{BoichenkoL-1998}
V.~Boichenko, G.~Leonov, Lyapunov's direct method in estimates of topological
  entropy, Journal of Mathematical Sciences 91~(6) (1998) 3370--3379.

\bibitem{PogromskyM-2011}
A.~Pogromsky, A.~Matveev, Estimation of topological entropy via the direct
  {L}yapunov method, Nonlinearity 24~(7) (2011) 1937--1959.

\end{thebibliography}

\documentclass[aps, amsmath, amssymb, aps, pre,twocolumn,groupedaddress,superscriptaddress,nofootinbib,longbibliography]{revtex4-1}

\usepackage{graphicx}
\usepackage[suffix=, outdir=./figs/]{epstopdf}
\usepackage[bottom]{footmisc}
\usepackage{enumitem}

\usepackage{array}
\usepackage{xfrac}
\usepackage[colorlinks,linkcolor=black,citecolor=blue,urlcolor=black]{hyperref}

\DeclareMathOperator{\LEs}{LE}

\newtheorem{theorem}{Theorem}
\newtheorem{lemma}{Lemma}
\newtheorem{corollary}{Corollary}

\begin{document}

\title{The Lyapunov dimension, convergency and entropy \\
for a dynamical model of Chua memristor circuit}

\author{G.A. Leonov}
\affiliation{Faculty of Mathematics and Mechanics, St. Petersburg State University,
Peterhof, St. Petersburg, Russia}
\author{N.V. Kuznetsov}
\email[]{Corresponding author: nkuznetsov239@gmail.com}
\affiliation{Faculty of Mathematics and Mechanics, St. Petersburg State University,
Peterhof, St. Petersburg, Russia}
\affiliation{Department of Mathematical Information Technology,
University of Jyv\"{a}skyl\"{a}, Jyv\"{a}skyl\"{a}, Finland}
\date{\today}

\keywords{Henon map, self-excited attractor, hidden attractor, chaos}

\begin{abstract}
\end{abstract}
\maketitle

\section*{Introduction}
For the study of chaotic dynamics and dimension of attractors
the concepts of the Lyapunov exponents \cite{Lyapunov-1892}
was found useful and became widely spread
\cite{GrassbergerP-1983,EckmannR-1985,ConstantinFT-1985,AbarbanelBST-1993}.
Such characteristics of chaotic behavior, as
the Lyapunov dimension \cite{KaplanY-1979}
and the entropy rate \cite{Kolmogorov-1959,Sinai-1959,AdlerKA-1965},
can be estimated via the Lyapunov exponents
\cite{Millionschikov-1976,Pesin-1977,KaplanY-1979,Young-2013}.
In this work an analytical approach to the study
of the Lyapunov dimension, convergency and entropy
for a dynamical model of Chua memristor circuit is demonstrated.

\section{A dynamical model of the Chua memristor circuit}

Consider one of the Chua memristor models \cite[eq.25]{CorintoF-2017}
\begin{equation}\label{chua-memristor}
\begin{aligned}
  & \dot x = \alpha(m_0-1)x+\alpha y-\alpha m_1x^3+\alpha x_0,  \\
  & \dot y = x-y+z, \ \dot z = \beta y-\gamma z
\end{aligned}
\end{equation}
with real parameters $\alpha,\beta,m_0,m_1,\gamma$, $x_0$,
and suppose that
\(
  \alpha m_1 > 0.
\)
For a survey on memristor circuits see, e.g. \cite{Tetzlaff-2014-book}.
System \eqref{chua-memristor} with $x_0=0$  describes the dynamics
of the Chua oscillator with cubic nonlinearity \cite{Chua-1993,HuangPWF-1996}.
If $\gamma=0$ and $x_0=0$, %$|x_0| >  \frac{2}{3}(m_0-1)\sqrt{\frac{m_0-1}{3m_1}}$
then system \eqref{sys:ode} has equilibria
$u^{eq}_0=(0,0,0)$ and
$u^{eq}_\pm = (\pm \sqrt{\frac{m0 - 1}{m_1}},0,\mp \sqrt{\frac{m0 - 1}{m_1}})$ for $m_0>1$.
Represent system \eqref{chua-memristor}
as an autonomous differential equation of general form:
\begin{equation}\label{sys:ode}
  \dot{u} = f({u}),
\end{equation}
where $u=(x,y,z) \in U=\mathbb{R}^3$ and the
continuously differentiable vector-function $f: \mathbb{R}^3 \to \mathbb{R}^3$
is the right-hand side of system \eqref{chua-memristor}.
Define by $u(t,u_0)$ a solution of \eqref{sys:ode} such that $u(0,u_0)=u_0$,
and consider the evolutionary operator $\varphi^t(u_0) = u(t,u_0)$.
We assume the uniqueness and existence of solutions of \eqref{sys:ode} for $t \in [0,+\infty)$.
Then system \eqref{sys:ode} generates a dynamical system
$\{\varphi^t\}_{t\geq0}$.
Let a nonempty set $K \subset U$ be invariant with respect to $\{\varphi^t\}_{t\geq0}$,
i.e. $\varphi^t(K) = K$ for all $t \geq 0$.
For example, as the set $K$ one can consider various types of attractors
of system \eqref{sys:ode} (see, e.g. examples from \cite{HuangPWF-1996,BilottaP-2008}).
Recently the \emph{classification of local attractors as being hidden or self-excited}
was introduced in connection with the discovery of the first hidden attractor
in the classical Chua model with the saturation nonlinearity
\cite{KuznetsovLV-2010-IFAC,LeonovKV-2011-PLA,BraginVKL-2011,LeonovKV-2012-PhysD,KuznetsovKLV-2013,KiselevaKKKLYY-2017,StankevichKLC-2017}:
an attractor is called a \emph{self-excited attractor}
if its basin of attraction
intersects with any open neighborhood of an equilibrium,
otherwise, it is called a \emph{hidden attractor} \cite{LeonovKV-2011-PLA,LeonovK-2013-IJBC,LeonovKM-2015-EPJST,Kuznetsov-2016}. %,
For example, hidden attractors can be found in various memristive circuits,
see, e.g. \cite{PhamVVJKH-2016,HuLLYZ-2017-HA,BaoBWCX-2017-HA,BaoBLWW-2016-HA,PhamVVHV-2016-HA,Vaidyanathan-2017-SCI-HA,VaidyanathanPV-2017-HA,PhamVVTT-2017-HA,PhamVVWH-2017-HA,RochaRK-2017-HA,SahaSRC-2015-HA,SemenovKASVA-2015,PhamVJWV-2014-HA,PhamVVLV-2015-HA,ChenYB-2015-HA-EL,ChenLYBXW-2015-HA}.
Remark that hidden attractors are not connected with equilibria and, thus,
do not related with the Shilnikov scenario of chaos \cite{AfraimovichGLShT-2014}.
In the works \citep{BianchiKLYY-2015,BlagovKLYY-2015,LeonovKYY-2015-TCAS,KuznetsovLYY-2017-CNSNS}
it is demonstrated the difficulties of reliable simulation
of the phase-locked loops circuits in SPICE and MATLAB Simulink,
caused by hidden attractors with narrow basins of attraction.

%For $x_0=0$ system \eqref{sys:lorenz-general}
%has the following equilibria: $u^{eq}_0 = (0,0,0)$ and $u^{eq}_{\pm}=$
%\[
%  \bigg
%  (\pm \tfrac{\sqrt{\beta m_0 - \beta + \gamma m_0}}{\sqrt{m_1(\beta + \gamma)}},
%  \pm \tfrac{\gamma\sqrt{\beta m_0 - \beta + \gamma m_0}}{\sqrt{m_1(\beta + \gamma)^3}},
%  \mp \tfrac{\beta\sqrt{\beta m_0 - \beta + \gamma m_0}}{\sqrt{m_1(\beta + \gamma)^3}}
%  \bigg)
%\]
%for $\beta m_0 - \beta + \gamma m_0>0$ and $\beta\sqrt{\beta m_0 - \beta + \gamma m_0>0$.

Consider linearization of system \eqref{sys:ode}
along the solution $u(t,u_0)=\varphi^t(u)$:
\begin{equation} \label{sfl}
  \begin{aligned}
    & \dot v = J(\varphi^t(u))v,
    \quad J(u) = Df(u),
  \end{aligned}
\end{equation}
where $J(u)$ is the $3\!\times\!3$ Jacobian matrix
\[
  J(u) =\left(
       \begin{array}{ccc}
         \alpha(m_0-1)-3\alpha m_1x^2 & \alpha & 0 \\
         1 & -1 & 1 \\
         0 & \beta & -\gamma \\
       \end{array}
     \right)
\]
and it can be represented as
\(
  J=J(0)-3\alpha m_1x^2I_1
\) with
\[
J_0= J(0) =\left(
       \begin{array}{ccc}
         \alpha(m_0-1)&\alpha & 0 \\
         1 & -1 & 1 \\
         0 & \beta & -\gamma \\
       \end{array}
     \right), \
I_1 =\left(
       \begin{array}{ccc}
         1&0 & 0 \\
         0 & 0 & 0 \\
         0 & 0 & 0 \\
       \end{array}
     \right).
\]

Let for any $t > 0$ and any $u \in U$ the ordered sequence
$\lambda_1(u) \ge\!\cdots\!\ge \lambda_n(u)$, where
$\lambda_i(u) = \lambda_i(\tfrac{1}{2} (J(u) + J(u)^{*})$,
$i = 1,\dots,n$ be the eigenvalues of the symmetrized Jacobian matrix
\(
%\begin{equation} \label{SJS}
  \tfrac{1}{2} \left(J(u) + J(u)^{*} \right).
%\end{equation}
\)
\begin{lemma}
\(
 \lambda_j(0) \geq \lambda_j(u), \quad j=1,2,3
\)
\end{lemma}

Then from Corollary~\ref{thm:dLKYeig} (see in the Appendix) we get

\begin{theorem}
\(
 d_{\rm L}^{\rm KY}\big(\{\lambda_{j}(u)\}_{i=1}^3\big)
 \leq
 d_{\rm L}^{\rm KY}\big(\{\lambda_{j}(0)\}_{i=1}^3\big)
\)
\end{theorem}

If $J(0)$ have simple real eigenvalues $\lambda_i(J(0))$, then
$\lambda_i(J(0)) = \lambda_i(\tfrac{1}{2}(J(0)+J^*(0)))$ and
we get the following result

\begin{corollary}
Let $u^{eq}_0=(0,0,0)$ be one of the equilibria of system \eqref{chua-memristor}
and the matrix $J(0)$ have simple real eigenvalues.
Then the exact Lyapunov dimension
of any compact invariant set $K \ni u^{eq}_0$ is defined as
\[
 \dim_{\rm L}K =
 d_{\rm L}^{\rm KY}\big(\{\lambda_{j}(0)\}_{i=1}^3\big).
\]
\end{corollary}

By Theorem~\ref{theorem:l1l2} (see in the Appendix) we get
\begin{theorem}
If
\(
 \lambda_{1}(0)+\lambda_{2}(0) <0,
\)
then any bounded solution of system \eqref{sys:ode}
tends to the stationary set.
\end{theorem}

%For example, consider \cite{CorintoF-2017}
%system \eqref{chua-memristor} with parameters
%$\beta = 16$, $m_0 = 1-0.143$, $m_1 = 1$, $\alpha=10$, and $x_0=0$.
%Then \eqref{chua-memristor} has equilibria
%$u^{eq}_0=(0,0,0)$ and
%$u^{eq}_\pm = (\pm 3/2,0, \mp 3/2)$.
%
%Let $m_1=1$, $c=1-m_0$, $\alpha=a$ and $\beta=-b$
%\[
%\begin{aligned}
%  & \dot x = -acx+ay-ax^3,  \\
%  & \dot y = x-y+z, \\
%  & \dot x = -by
%\end{aligned}
%\]
%Characteristic polynomial is
%\[
%  t = {\rm Tr}(J(u)) = - 3ax^2 - ac - 1
%\]
%\[
%  \det(pI-J(u)) =
%    p^3 + (3ax^2 + ac + 1)p^2 + (3ax^2 - a + b + ac)p + 3abx^2 + abc
%\]
%\[
%  \det(pI-J(u)) =p^3 -t p^2 + (-t-1-a+b)p + b(-t-1)
%\]
%\[
%p_1(x)+p_2(x)+p_3(x) = t = - 3ax^2 - ac - 1
%\]

\newpage
\section*{Appendix. Exact and finite-time Lyapunov dimension}
Suppose that $\det J(u) \neq 0 \quad \forall u \in U$.
Consider a fundamental matrix of linearized system \eqref{sfl}
$D\varphi^t(u)$ such that $D\varphi^0(u) = I$, where $I$ is a unit $3 \times 3$ matrix.
Let $\sigma_i(t,u) = \sigma_i(D\varphi^t(u))$, $i = 1,2,3$,
be the singular values of $D\varphi^t(u)$
with respect to their algebraic multiplicity
ordered so that $\sigma_1(t,u) \geq \sigma_2(t,u) \geq \sigma_3(t,u) > 0$
for any $u \in U$ and $t \geq 0$.
Consider the set of \emph{the finite-time Lyapunov exponents} at the point $u_0$ $\{\LEs_i(t,u_0)=\frac{1}{t}\ln\sigma_{j}(t,\!u_0)\}_{i=1}^3$
ordered by decreasing  for $t > 0$.

Introduce the \emph{Kaplan-Yorke formula \cite{KaplanY-1979} with respect
to the ordered set} $\lambda_1\geq... \geq \lambda_n$:
\begin{equation}\label{lftKY}
   d^{\rm KY}(\{\lambda_i\}_{i=1}^3)\!=\!
   j+\frac{\sum_{i=1}^{j}\lambda_i}{|\lambda{j + 1}|},  \
   j\!=\!\max\{m\!: \sum_{i=1}^{m}\!\lambda_i\!\geq\!0\},
\end{equation}
where $d^{\rm KY}(\{\lambda_i\}_{i=1}^3)=0$ for $j=0$ and
$d^{\rm KY}(\{\lambda_i\}_{i=1}^3)=3$ for $j=3$.
Then
the \emph{finite-time local Lyapunov dimension}
\cite{Kuznetsov-2016-PLA,KuznetsovLMPS-2017}
at a certain point $u_0$ can be defined as
\[
   \dim_{\rm L}(t,u_0) = d^{\rm KY}(\{\LEs_i(t,u_0)\}_{i=1}^3).
\]
and the \emph{finite-time Lyapunov dimension}
of invariant closed bounded set $K$
is as follows
\begin{equation}\label{DOmaptmax}
  \dim_{\rm L}(t, K) = \sup\limits_{u_0 \in K} \dim_{\rm L}(t,u_0).
\end{equation}
In this approach the use of Kaplan-Yorke formula \eqref{lftKY} with
the finite-time Lyapunov exponents is justified
by the \emph{Douady--Oesterl\'{e} theorem} \cite{DouadyO-1980},
which implies that for any fixed $t > 0$
the Lyapunov dimension of the map $\varphi^t$ with respect
to a closed bounded invariant set $K$, defined by \eqref{DOmaptmax},
is an upper estimate of the Hausdorff dimension of the set $K$:
\(
  \dim_{\rm H}K \leq \dim_{\rm L}(t, K).
\)
For the estimation of the Hausdorff dimension of invariant closed bounded set $K$
one can use the map $\varphi^t$ with any time $t$
(e.g. $t=0$ leads to the trivial estimate $\dim_{\rm H}K \leq 3$)
and, thus, the best estimation is
\(
  \dim_{\rm H}{K} \le \inf_{t\geq0}\dim_{\rm L} (t, K).
\)
The following property
\begin{equation}\label{DOlim}
  \inf_{t\geq0}\sup\limits_{u \in K} \dim_{\rm L}(t,u)
  = \liminf_{t \to +\infty}\sup\limits_{u \in K} \dim_{\rm L}(t,u)
\end{equation}
allows one to introduce the \emph{Lyapunov dimension}  %of $K$ as
\cite{Kuznetsov-2016-PLA}
\begin{equation}\label{DOinf}
  \dim_{\rm L} K
  = \liminf_{t \to +\infty}\sup\limits_{u \in K} \dim_{\rm L}(t,u).
\end{equation}
If the maximum of local Lyapunov dimensions on the global attractors,
which involves all equilibria, is achieved at an equilibrium point $u^{cr}_{eq}$,
i.e. $\dim_{\rm L} u^{cr}_{eq} = \max_{u_0 \in K} \dim_{\rm L} u_0$,
then this allows one to get the \emph{exact Lyapunov dimension}.
(this term was suggested by Doering~et~al.~in~\cite{DoeringGHN-1987}).
In general, a \emph{conjecture on the Lyapunov dimension of self-excited attractor} \cite{KuznetsovLMPS-2017}
is that for a typical system
the Lyapunov dimension of a self-excited attractor
does not exceed the Lyapunov dimension of one of unstable equilibria,
the unstable manifold of which intersects with the basin of attraction
and visualize the attractor.

%Recall that a set $K$ with noninteger Hausdorff dimension
%is referred to as a \emph{fractal set} \cite{EckmannR-1985}
%and, when such set $K$ is an attractor,
%it is called a \emph{strange attractor}
%\cite{RuelleT-1971,GrebogiOY-1987}.

In contrast to the finite-time Lyapunov dimension \eqref{DOmaptmax},
the Lyapunov dimension \eqref{DOinf}
is \emph{invariant under smooth change of coordinates}
\cite{KuznetsovAL-2016,Kuznetsov-2016-PLA}.
This property and a proper choice of smooth change of coordinates
may significantly simplify the estimation of the Lyapunov dimension
of dynamical system.

%Based on this property and using a proper choice of smooth change of coordinates,
%in many cases it is possible to compute the exact Lyapunov dimension analytically
%by the method,  prosed by Leonov \cite{Leonov-1991-Vest,LeonovB-1992,Leonov-2012-PMM,Kuznetsov-2016-PLA}.

Consider an effective analytical approach, proposed by Leonov
\cite{Leonov-1991-Vest,LeonovB-1992,Kuznetsov-2016-PLA}.
Let for any $t > 0$ and any $u_0 \in U$ the ordered sequence
$\lambda_1(u_0, S) \ge\!\cdots\!\ge \lambda_n(u_0, S)$, where
$\lambda_i(u_0, S) = \lambda_i(S \varphi^t(u_0))$, $i = 1,\dots,n$
be the eigenvalues of the symmetrized Jacobian matrix
\begin{equation} \label{SJS}
  \frac{1}{2} \left( S J(\varphi^t(u_0)) S^{-1} + (S J(\varphi^t(u_0)) S^{-1})^{*}\right).
\end{equation}

\begin{theorem}\label{theorem:th1}
If there exist an integer $j \in \{1,\ldots,n-1\}$,
a real $s \in [0,1]$, a differentiable scalar function $V: U \subseteq \mathbb{R}^n \to \mathbb{R}^1$, and a nonsingular $n\times n$ matrix $S$
such that
\begin{equation}\label{ineq:weilSVct}
  \sup_{u \in U} \big( \lambda_1 (u,S) + \cdots + \lambda_j (u,S)
  + s\lambda_{j+1}(u,S) + \dot{V}(u) \big) < 0,
\end{equation}
where $\dot{V} (u) = ({\rm grad}(V))^{*}f(u)$,
then for a compact invariant set $K\subset U$ we have
\[
   \dim_{\rm H}K \leq
    \dim_{\rm L}(\{\varphi^t\}_{t\geq0},K)
   < j+s.
\]
\end{theorem}
%This method is based on a combination of the Douady-Oesterl\'{e} approach
%with the direct Lyapunov method and
In the work \cite{Kuznetsov-2016-PLA}
it is shown how the method can be justified
by the invariance of the Lyapunov dimension of compact invariant set
with respect to the special smooth change of variables $h$
with $Dh(u)=e^{V(u)(j+s)^{-1}}S$, where $V$
is a differentiable scalar function and $S$ is a nonsingular $n \times n$ matrix.
For $S=0$ and $V(u) \equiv 0$ we have
\begin{corollary}[\cite{DouadyO-1980,Smith-1986,Leonov-1991-Vest,Kuznetsov-2016-PLA}]
\label{thm:dLKYeig}
%For a compact invariant set $K$ %we have
%\begin{equation}\label{dLKYeig}
\[
 \dim_{\rm H}K\!\leq\!\dim_{\rm L}K \!\leq\!
 \sup_{u \in K}d_{\rm L}^{\rm KY}\!\big(\{\lambda_{j}(u)\}_{i=1}^3\big).
\]
\end{corollary}

The following result~\cite{Leonov-2012-PMM} is useful for the study of global convergency.
\begin{theorem}\label{theorem:l1l2}
If there exist a continuously differentiable scalar function
$V: U \subseteq \mathbb{R}^n \to \mathbb{R}^1$ and
a non-degenerate $n\times n$ matrix $S$ exist such that
\begin{equation}\label{pmm36}
  \sup_{u \in U} \big(
  \lambda_1(u,S)+\lambda_2(u,S)+\dot V(u)
  \big)
  <0,
\end{equation}
then any bounded solution of system \eqref{sys:ode}
with any initial data $u_0 \in U$
tends to the stationary set of dynamical system $\{\varphi^t\}_{t\geq0}$ as $t\to+\infty$.
\end{theorem}
Remark that the stationary set can have any structure,
e.g. %can consists of a finite number of stable and unstable point or
be a line of  equilibria.

In \cite{BoichenkoL-1998,PogromskyM-2011} it is demonstrated how the above technique
can be effectively used to a derive constructive upper bound of
the sum of positive Lyapunov exponents and
the topological entropy \cite{AdlerKA-1965}
(the topological entropy is an analogue of the entropy defined earlier
by Kolmogorov and Sinai \cite{Kolmogorov-1959,Sinai-1959}).

\section*{Acknowledgment}
\vspace{-0.5cm}
We would like to thank Leon Chua, Fernando Corinto, and
Ronald Tetzlaff to draw our attention to the memristors
and for fruitful discussions.

\vspace{-0.5cm}
%\bibliographystyle{elsarticle-num}
%\bibliography{C:/Dropbox/bib/bib_nk,C:/Dropbox/bib/bib_leonov,C:/Dropbox/bib/bib_full,C:/Dropbox/bib/genlorenz-bib,C:/Dropbox/bib/bib-hidden}

\end{document}